OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations

OpenStudy (anonymous):

x + 5y = 3 2x + 9y = 4

OpenStudy (anonymous):

OMG this is really over the top.

OpenStudy (anonymous):

x = y+ 40(5y) x = 56 + e^y x = -7 therefore y = 2